# On polynomial interpolation in the monomial basis

### v4

@@ 2.3 Interpolation over an interval
Figure 12: Polynomial interpolation in the monomial basis over $\Gamma=[0,1]$. The constant $\rho_{*}$ equals $3+2\sqrt{2}$.
2.4 Interpolation over a smooth simple arc in the complex plane
In this section, we consider polynomial interpolation in the monomial basis over a smooth simple arc $\Gamma\subset\mathbbm{C}$. In this more general setting, similar to the special case where $\Gamma$ is an interval, there exists a class of collocation points, known as adjusted Fejér points, whose associated Lebesgue constant also grows logarithmically [29]. However, these points are extremely costly to construct numerically. On the other hand, the set of collocation points constructed based on the following procedure, while suboptimal, is a good choice for practical applications. Suppose that $g:[-1,1]\to\mathbbm{C}$ is a parameterization of $\Gamma$. Provided that the Jacobian $g^{\prime}(t)$ does not have large variations, we find that the Lebesgue constant for the set of collocation points $Z=\{g(t_{j})\}_{j=0,1,\dots,N}$, where $\{t_{j}\}_{j=0,1,\dots,N}$ is the set of $(N+1)$ Chebyshev points on the interval $[-1,1]$, grows at a slow rate. It is worth noting that $\{t_{j}\}_{j=0,1,\dots,N}$ can also be chosen as the Legendre points on the interval $[-1,1]$, for the same reason stated in Remark 2.8.
In the rest of this section, we provide several numerical experiments involving interpolation over smooth simple arcs in the complex plane. In particular, we consider the scenario where $\Gamma$ is a parabola parameterized by $g:[-1,1]\to\mathbbm{C}$, $g(t):=t+i\alpha(t^{2}-1)$, for $\alpha=0.2$, $0.4$, $0.6$. In Figure 13, we plot these parabolas, including their associated level sets $E_{\rho}$, for various values of $\rho$. The value of $\rho_{*}$ for each parabola is estimated from the plots. In Figure 14, we estimate the condition numbers of the Vandermonde matrices and the Lebesgue constants for the sets of collocation points, for different values of $\alpha$. One can observe that the Lebesgue constants are of approximately size one, which justifies our choice of collocation points. In Figure 15, we report the monomial approximation error ${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Gamma)}$, and the estimated values of $u\cdot{\lVert a^{(N)}\rVert}_{2}$, for various functions $F$ over $\Gamma$. Based on the experimental results, it is clear that the observations made at the end of Section 2.3 are also applicable to the case where $\Gamma$ is a parabola. In fact, these observations apply to any simple arc that is sufficiently smooth.
(a) $g(t)=t+i\alpha(t^{2}-1)$
(b) $\alpha=0.2,\,\rho_{*}\approx 2.56$
(c) $\alpha=0.4,\,\rho_{*}\approx 2.6$
(d) $\alpha=0.6,\,\rho_{*}\approx 2.6$
Figure 13: The level set $E_{\rho}$ of a parabola, for various values of $\rho$. The colorbar indicates the value of $\rho$. The smooth simple arc $\Gamma$ is the white curve in the figure. The value of $\rho_{*}$ (see Definition 2.2) is estimated for each arc $\Gamma$.
Figure 14: The Lebesgue constant for collocation points over a parabola, and the 2-norm of the inverse of the corresponding Vandermonde matrix. The collocation points are chosen to be $\{g(t_{j})\}$, where $\{t_{j}\}$ is a set of Chebyshev points over $[-1,1]$, and $g:[-1,1]\to\mathbbm{C}$ is the parameterization of the parabola defined in Section 2.4. The $x$-axis label $N$ denotes the order of approximation. The value of $\rho_{*}$ is set to be $2.6$, based on the estimate in Figure 13.
(a) $F(z)=\cos(2z+1)$
(b) $F(z)=\cos(8z+1)$
(c) $F(z)=e^{-z^{2}}$
(d) $F(z)=\frac{1}{z-2}$
(e) $F(z)=\frac{1}{z+i}$
(f) $F(z)=\tan(\tan(z)/2)$
Figure 15: Polynomial interpolation in the monomial basis over a parabola. The interpolation is performed on the parabolas shown in Figure 12(a).
@@ 2.4 Interpolation over a smooth simple arc in the complex plane
In certain applications, the function $F:\Gamma\to\mathbbm{C}$ is defined by the formula $F(z):=\sigma(g^{-1}(z))$, where $g:[-1,1]\to\mathbbm{C}$ is an analytic function that parameterizes the curve $\Gamma$, and $\sigma:[-1,1]\to\mathbbm{C}$ is analytic. In this case, the analytic continuation of $F$ can have a singularity close to $\Gamma$ even when $\sigma$ is entire, because the inverse of the parameterization (i.e., $g^{-1}$) has so-called Schwarz singularities at $z=g(t^{*})$, where $g^{\prime}(t^{*})=0$. In [2], the authors show that, the higher the curvature of the arc $\Gamma$, the closer the singularity induced by $g^{-1}$ is to $\Gamma$. As a result, the approximation of such a function $F$ by polynomials is efficient only when the curvature of $\Gamma$ is small.
@@ 3 Applications
After justifying the use of a monomial basis for polynomial interpolation, a natural question to ask is: why would one want to do it in the first place? For one, the monomial basis is the simplest polynomial basis to manipulate. For example, the evaluation of an $N$th degree polynomial expressed in the monomial basis can be achieved using only $N$ multiplications through the application of Horner’s rule. This evaluation can be further accelerated using Estrin’s scheme, which has distinct advantages on modern processors. Additionally, the derivative and anti-derivative of an $N$th degree polynomial in the monomial basis can be calculated more stably in other bases, and using only $N$ multiplications. Besides these obvious advantages, we discuss some other applications below.
@@ 3.1 Oscillatory integrals and singular integrals
Given an oscillatory (or singular) function $\Psi:\Gamma\to\mathbbm{C}$ and a smooth function $F:\Gamma\to\mathbbm{C}$ over a smooth simple arc $\Gamma\subset\mathbbm{C}$, the calculation of $\int_{\Gamma}\Psi(z)F(z)\,\mathrm{d}z$ by standard quadrature rules can be extremely expensive or inaccurate due to the oscillations (or the singularity) of $\Psi$. However, when $F$ is a monomial, there exists a wide range of integrals in the form above that can be efficiently computed to high accuracy by either analytical formulas or by recurrence relations, often derived using integration by parts. Therefore, when the smooth function $F$ is accurately approximated by a monomial expansion of order $N$, such integrals can be efficiently evaluated by the formula $\sum_{k=0}^{N}a_{k}\Bigl{(}\int_{\Gamma}\Psi(z)z^{k}\,\mathrm{d}z\Bigr{)}$, where $\{a_{k}\}_{k=0,1,\dots,N}$ denotes the coefficients of the monomial expansion. Integrals of this type include the Fourier integral $\int_{a}^{b}e^{icx}F(x)\,\mathrm{d}x$, and various layer potentials, e.g., $\int_{\Gamma}\log(z-\xi)F(z)\,\mathrm{d}z$ and $\int_{\Gamma}\frac{F(z)}{z-\xi}\,\mathrm{d}z$, where $\xi\in\mathbbm{C}$ is given. We refer the readers to [20, 21] for more detailed discussion on the Fourier integral, and to [17, 2, 24] for more detailed discussion on the application of polynomial interpolation in the monomial basis to the evaluation of layer potentials. Some interesting applications can also be found in [18, 1, 22].
@@ 3.2 Root finding
Given a smooth simple arc $\Gamma\subset\mathbbm{C}$ and a function $F:\Gamma\to\mathbbm{C}$, one method for computing the roots of $F$ over $\Gamma$ is to first approximate it by a polynomial $P_{N}(z)=\sum_{j=0}^{N}a_{j}z^{j}$ to high accuracy, and then to compute the roots of $P_{N}$ by calculating the eigenvalues of the corresponding companion matrix. Recently, a backward stable algorithm that computes the eigenvalues of $C(P_{N})$ in $\mathcal{O}(N^{2})$ operations with $\mathcal{O}(N)$ storage has been proposed in [5]. This algorithm is backward stable in the sense that the computed roots are the exact roots of a perturbed polynomial $\widehat{P}{}_{N}(z)=\sum_{j=0}^{N}(a_{j}+\delta a_{j})z^{j}$, so that the backward error satisfies ${\lVert\delta a^{(N)}\rVert}_{2}\lesssim u{\lVert a^{(N)}\rVert}_{2}$, where $u$ denotes machine epsilon, $\delta a^{(N)}:=(\delta a_{0},\delta a_{1},\dots,\delta a_{N})^{T}$ and $a^{(N)}:=(a_{0},a_{1},\dots,a_{N})^{T}$. It follows that ${\lVert P_{N}-\widehat{P}_{N}\rVert}_{L^{\infty}(\Gamma)}\leq u{\lVert\delta a^{(N)}\rVert}_{1}\lesssim u\sqrt{N+1}{\lVert a^{(N)}\rVert}_{2}$. When ${\lVert a^{(N)}\rVert}_{2}\approx{\lVert P_{N}\rVert}_{L^{\infty}(\Gamma)}$, the computed roots are backward stable in the polynomial $P_{N}$. This condition, however, does not hold for all polynomials $P_{N}$. Furthermore, the calculation of the coefficients $a^{(N)}$ from the function $F$, which involves the solution of a Vandermonde system of equations, is highly ill-conditioned. In this paper, we show that, when $F$ is sufficiently smooth, it is possible to compute the coefficients of an interpolating polynomial $P_{N}(z)=\sum_{j=0}^{N}a_{j}z^{j}$, with ${\lVert a^{(N)}\rVert}_{2}\approx{\lVert F\rVert}_{L^{\infty}(\Gamma)}$, which approximates $F$ uniformly to high accuracy, even when the condition number of the Vandermonde matrix is close to the reciprocal of machine epsilon. From this, we see that a backward stable root finder can be constructed by combining the piecewise polynomial approximation procedure described in Section 2.2 with the algorithm presented in [5],
@@ 4 Discussion
Since the invention of digital computers, most research on the topic of polynomial interpolation in the monomial basis focuses on showing that it is a bad idea. The condition number of Vandermonde matrices has been studied extensively in recent decades (see [14] for a literature review), and it is known that its growth rate is at least exponential, unless the collocation nodes are distributed uniformly on the unit circle centered at the origin [23]. As a result, the computed monomial coefficients are generally highly inaccurate when the dimensionality of the Vandermonde matrix is not small. For this reason, other more well-conditioned bases are often used for polynomial interpolation [27, 11]. On the other hand, it has long been observed that polynomial interpolation in the monomial basis produces highly accurate approximations for sufficiently smooth functions (see, for example, [16, 17]). This is because that the inaccurately computed monomial coefficients does not imply that resulting interpolating polynomial is bad, since it is the backward error ${\lVert V\widehat{a}-f\rVert}_{2}$ of the numerical solution $\widehat{a}$ to the Vandermonde system $Va=f$ that determines the accuracy of the approximation, and ${\lVert V\widehat{a}-f\rVert}_{2}$ can be small even when the condition number $\kappa(V)$ is large. It has been shown in both frame approximations [3, 4] and the method of fundamental solutions [6, 26] that ${\lVert V\widehat{a}-f\rVert}_{2}\lesssim u\cdot{\lVert a\rVert}_{2}$, where $u$ denotes machine epsilon, from which it is easy to derive that the monomial approximation error is bounded by the sum of the polynomial interpolation error and the extra error term $u\cdot{\lVert a\rVert}_{2}$. In this paper, we characterize the growth of ${\lVert a\rVert}_{2}$, and show that this extra error term is generally smaller than the polynomial interpolation error, provided that the order of approximation is no larger than the maximum order allowed by the constraint $\kappa(V)\lesssim\frac{1}{u}$. Since this maximum order is not small in practice, we find that the monomial basis is a useful basis for interpolation, especially when it is used to construct a piecewise polynomial approximation.
@@
While not discussed in this paper, our results can be easily generalized to higher dimensions. In [25], we study bivariate polynomial interpolation in the monomial basis over a (possibly curved) triangle, and demonstrate that the resulting order of approximation can reach up to 20, regardless of the triangle’s aspect ratio.

### v5

@@ 3.1 Interpolation over an interval
Figure 12: Polynomial interpolation in the monomial basis over $\Omega=[0,1]$. The constant $\rho_{*}$ equals $3+2\sqrt{2}$. The value of $C$ is chosen such that ${\lVert F-P_{N}\rVert}_{L^{\infty}([0,1])}\leq C\rho_{*}^{-N}$ for $N=0,1,...,20$. We highlight the region bounded above by $u\cdot C$ in pink.
(a) $F(x)=e^{-2(x+0.1)^{2}}$
(b) $F(x)=\cos(12x+1)$
(c) $F(x)={\lvert x+0.1\rvert}^{2.5}$
(d) $F(x)=T_{30}(x)$
Figure 13: Polynomial interpolation in the monomial basis over $\Omega=[-1,1]$, when the order of approximation exceeds the threshold. The vertical line with the label ‘Threshold’ denotes the order of approximation $\min\bigl{(}\{N:{\lVert(V^{(N)})^{-1}\rVert}_{2}\geq\frac{1}{u}\}\bigr{)}$, beyond which our theory is no longer applicable.
3.2 Approximation over more general regions in the complex plane
In this section, we consider polynomial interpolation and approximation in the monomial basis over a more general simply connected compact set $\Omega\subset\mathbbm{C}$. When $\partial\Omega$ is an analytic Jordan curve, it is shown in [28] that the Lebesgue constant of the Fejér points $\{\Phi^{-1}\bigl{(}\exp(i(\frac{2\pi j}{N+1}+a))\bigr{)}\}_{j=0,1,\dots,N}$ grows logarithmically, where $\Phi^{-1}$ is defined in Remark 2.3, and where $a\in\mathbbm{R}$ is arbitrary. When $\Omega$ is a Jordan arc, it is shown in [38] that the Lebesgue constant of the Fejér points, with some adjustment to ensure adequate spacing between interpolation points, also exhibits logarithmic growth.
In the remainder of this section, we provide several numerical experiments involving the following three domains: an ellipse with a major radius of 1 and a minor radius of $0.2$, centered at the origin (see Figure 6(d)); a parabola parameterized by $g:[-1,1]\to\mathbbm{C}$, where $g(t):=t+0.4i(t^{2}-1)$ (see Figure 6(e)); and a square with side length $\sqrt{2}$, centered at the origin. We note that the estimated constants $\rho_{*}$ (see Definition 2.2) for the three domains are approximately $2$, $2.6$, and $1.3$, respectively (see Figures 6(d), 6(e), and 6(f)).
In the first domain (the ellipse), the conformal mapping $\Phi^{-1}$ is specified by the formula $\Phi^{-1}(z)=\frac{3}{5}\cdot z+\frac{2}{5}\cdot\frac{1}{z}$, and we use the Fejér points for interpolation. Since the Fejér points for the second and the third domains are not known in closed form, we employ different approaches tailored to each case. In the second domain (the parabola), we set the interpolation points to be $\{g(t_{j})\}_{j=0,1,\dots,N}$, where $\{t_{j}\}_{j=0,1,\dots,N}$ is the set of $(N+1)$ Chebyshev points on the interval $[-1,1]$. We find that the Lebesgue constants associated with these points are all less than 10 when $\kappa(V^{(N)})\lesssim\frac{1}{u}$. In the third domain (the square), we instead solve a least-squares fitting problem with sampling points chosen to be the $2(N+1)$ Chebyshev points of the first kind along each side of the square, where $N$ denotes the order of approximation (see Remark 2.2).
In Figure 14, we report the values of ${\lVert(V^{(N)})^{-1}\rVert}_{2}$ and its upper bound $\rho_{*}^{N}\Lambda_{N}$ (see Theorem 2.5) for the ellipse domain and the parabola domain, where the value of $\rho_{*}$ for each domain is estimated based on the plots in Figure 6. Again, one can see that our upper bound is very tight.
(a) The ellipse, $\rho_{*}\approx 2$
(b) The parabola, $\rho_{*}\approx 2.6$
Figure 14: The 2-norm of $(V^{(N)})^{-1}$ and its upper bound.
In Figures 15, 16, and 17, we report the estimated values of ${\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}$, ${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Omega)}$, and $u\cdot{\lVert a^{(N)}\rVert}_{2}$, over each of these three domains. We estimate the exact polynomial interpolant $P_{N}$ by expressing $P_{N}$ in a discrete orthogonal polynomial basis (which is well-conditioned), computed using the Vandermonde with Arnoldi orthogonalization procedure [11]. In addition, we report the the upper bound $C\rho_{*}^{-N}$ for ${\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}$, where $C$ is chosen so that ${\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}\leq C\rho_{*}^{-N}$ for a range of $N$ indicated in the figure caption, and the upper bound $u\cdot C$ for $u\cdot{\lVert a^{(N)}\rVert}_{2}$. These results demonstrate that the observations made at the near end of Section 3.1 are indeed applicable to more general domains in the complex plane, and to polynomial approximation using least-squares fitting with monomials.
(a) $F(z)=\frac{1}{z-1.1}$, $C=1\times 10^{11}$
(b) $F(z)=e^{-4z^{2}}$, $C=3\times 10^{4}$
(c) $F(z)=\cos(12z+1)$, $C=2\times 10^{1}$
(d) $F(z)=\tan(\tan(z)/2)$, $C=5\times 10^{3}$
Figure 15: Polynomial interpolation in the monomial basis over the ellipse $\Omega$. We estimate $P_{N}$ by expressing it in a discrete orthogonal polynomial basis, computed using the Vandermonde with Arnoldi orthogonalization procedure [11]. The constant $\rho_{*}$ is approximately equal to $2.0$. The value of $C$ is chosen such that ${\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}\leq C\rho_{*}^{-N}$ for $N=0,1,...,55$. We highlight the region bounded above by $u\cdot C$ in pink.
(a) $F(z)=\frac{1}{z-1.1}$, $C=1\times 10^{11}$
(b) $F(z)=e^{-4z^{2}}$, $C=3\times 10^{6}$
(c) $F(z)=\cos(12z+1)$, $C=5\times 10^{2}$
(d) $F(z)=\tan(\tan(z)/2)$, $C=1\times 10^{5}$
Figure 16: Polynomial interpolation in the monomial basis over the parabola $\Omega$. We estimate $P_{N}$ by expressing it in a discrete orthogonal polynomial basis, computed using the Vandermonde with Arnoldi orthogonalization procedure [11]. The constant $\rho_{*}$ is approximately equal to $2.6$. The value of $C$ is chosen such that ${\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}\leq C\rho_{*}^{-N}$ for $N=0,1,...,40$. We highlight the region bounded above by $u\cdot C$ in pink.
(a) $F(z)=e^{-10z^{2}}$, $C=7\times 10^{3}$
(b) $F(z)=\frac{1}{z-0.8}$, $C=5\times 10^{4}$
(c) $F(z)=\cos(12z+1)$, $C=1\times 10^{9}$
(d) $F(z)=\tan(\tan(z))$, $C=5\times 10^{0}$
Figure 17: Polynomial approximation in the monomial basis over the square $\Omega$. We estimate $P_{N}$ by expressing it in a discrete orthogonal polynomial basis, computed using the Vandermonde with Arnoldi orthogonalization procedure [11]. The constant $\rho_{*}$ is approximately equal to $1.3$ (see Figure 6(f)). The value of $C$ is chosen such that ${\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}\leq C\rho_{*}^{-N}$ for $N=0,1,...,100$. We highlight the region bounded above by $u\cdot C$ in pink.
@@ 3.2 Approximation over more general regions in the complex plane
In certain applications [18, 2, 29], the function $F:\Gamma\to\mathbbm{C}$ is defined by the formula $F(z):=\sigma(g^{-1}(z))$, where $\Gamma$ is a simple smooth arc, $g:[-1,1]\to\mathbbm{C}$ is an analytic function that parameterizes $\Gamma$, and $\sigma:[-1,1]\to\mathbbm{C}$ is analytic. In this case, the analytic continuation of $F$ can have a singularity close to $\Gamma$ even when $\sigma$ is entire, because the inverse of the parameterization (i.e., $g^{-1}$) has so-called Schwarz singularities at $z=g(t^{*})$, where $g^{\prime}(t^{*})=0$. In [2], the authors show that, the higher the curvature of the arc $\Gamma$, the closer the singularity induced by $g^{-1}$ is to $\Gamma$. As a result, the approximation of such a function $F$ by polynomials is efficient only when the curvature of $\Gamma$ is small.
@@ 4 Applications
After justifying the use of a monomial basis for polynomial interpolation, a natural question to ask is why one would want to use it in the first place. For one, the monomial basis is the simplest polynomial basis to manipulate. For example, the evaluation of an $N$th degree polynomial expressed in the monomial basis can be achieved using only $N$ multiplications through the application of Horner’s rule. This evaluation can be further accelerated using Estrin’s scheme [13], which leverages instruction-level parallelism in modern CPUs. This type of parallelism is typically unavailable when evaluating a polynomial expressed in an orthogonal polynomial basis.
Besides this obvious advantage, we discuss some other applications below.
@@ 4.1 Oscillatory integrals and singular integrals
Given an oscillatory or singular function $\Psi:\Gamma\to\mathbbm{C}$ and a smooth function $F:\Gamma\to\mathbbm{C}$ over a smooth simple arc $\Gamma\subset\mathbbm{C}$, the calculation of $\int_{\Gamma}\Psi(z)F(z)\,\mathrm{d}z$ by standard quadrature rules can be extremely expensive or inaccurate due to the oscillations or the singularity of $\Psi$. However, when $F$ is a monomial, there exists a wide range of integrals in this form that can be efficiently computed to high accuracy by either analytical formulas or by recurrence relations, often derived using integration by parts. Therefore, when the smooth function $F$ is accurately approximated by a monomial expansion of order $N$, such integrals can be efficiently evaluated by the formula $\sum_{k=0}^{N}a_{k}\Bigl{(}\int_{\Gamma}\Psi(z)z^{k}\,\mathrm{d}z\Bigr{)}$, where $\{a_{k}\}_{k=0,1,\dots,N}$ denotes the coefficients of the monomial expansion. Integrals of this type include the Fourier integral $\int_{a}^{b}e^{icx}F(x)\,\mathrm{d}x$ and various layer potentials and Stieltjes transforms, e.g., $\int_{\Gamma}\log(z-\xi)F(z)\,\mathrm{d}z$ and $\int_{\Gamma}\frac{F(z)}{z-\xi}\,\mathrm{d}z$, where $\xi\in\mathbbm{C}$ is given. We refer the readers to [23, 24] for more detailed discussion on the Fourier integral, and to [18, 2, 29] for more detailed discussion on the application of polynomial interpolation in the monomial basis to the evaluation of layer potentials. Some interesting applications can also be found in [19, 1, 25].
@@ 4.2 Root finding
Given a simply connected compact set $\Omega\subset\mathbbm{C}$ and a function $F:\Omega\to\mathbbm{C}$, one method for computing the roots of $F$ over $\Omega$ is to first approximate $F$ by a polynomial $P_{N}(z)=\sum_{j=0}^{N}a_{j}z^{j}$ to high accuracy, and then to compute the roots of $P_{N}$ by calculating the eigenvalues of the corresponding companion matrix, which we denote by $C(P_{N})$. Recently, a backward stable algorithm that computes the eigenvalues of $C(P_{N})$ in $\mathcal{O}(N^{2})$ operations with $\mathcal{O}(N)$ storage has been proposed in [5]. This algorithm is backward stable in the sense that the computed roots are the exact roots of a perturbed polynomial $\widehat{P}_{N}(z)=\sum_{j=0}^{N}(a_{j}+\delta a_{j})z^{j}$, so that the backward error satisfies ${\lVert\delta a^{(N)}\rVert}_{2}\lesssim u{\lVert a^{(N)}\rVert}_{2}$, where $u$ denotes machine epsilon, $\delta a^{(N)}:=(\delta a_{0},\delta a_{1},\dots,\delta a_{N})^{T}$ and $a^{(N)}:=(a_{0},a_{1},\dots,a_{N})^{T}$. It follows that ${\lVert P_{N}-\widehat{P}_{N}\rVert}_{L^{\infty}(\Omega)}\leq{\lVert\delta a^{(N)}\rVert}_{1}\lesssim u\sqrt{N+1}{\lVert a^{(N)}\rVert}_{2}$. When ${\lVert a^{(N)}\rVert}_{2}\approx{\lVert P_{N}\rVert}_{L^{\infty}(\Omega)}$, the computed roots are backward stable in the polynomial $P_{N}$. This condition, however, does not hold for all polynomials $P_{N}$. Furthermore, the calculation of the coefficients $a^{(N)}$ from the function $F$, which involves the solution of a Vandermonde system of equations, is highly ill-conditioned. In this paper, we show that, when $F$ is sufficiently smooth (in the mild sense discussed in Section 2.1), it is possible to compute the coefficients of an interpolating polynomial $P_{N}(z)=\sum_{j=0}^{N}a_{j}z^{j}$, with ${\lVert a^{(N)}\rVert}_{2}\approx{\lVert F\rVert}_{L^{\infty}(\Omega)}$, which approximates $F$ uniformly to high accuracy, even when the condition number of the Vandermonde matrix is close to the reciprocal of machine epsilon. From this, we see that a backward stable root finder can be constructed by combining the piecewise polynomial approximation procedure described in Section 2.2 with the algorithm presented in [5]. Interestingly, by using the least-squares fitting procedure described in Section 3.2 for the case when $\Omega$ is a square, an adaptive root-finding algorithm for a complex analytic function in a square domain, analogous to the one described in [36], can be developed utilizing solely the monomial basis.
@@ 5 Discussion
Since the invention of digital computers, most research on the topic of polynomial interpolation in the monomial basis has focused on showing that it is a bad idea. The condition number of Vandermonde matrices has been studied extensively in recent decades (see [15] for a literature review), and it is known that its growth rate is at least exponential, unless the interpolation points are asymptotically distributed uniformly on the complex unit circle centered at the origin [26]. As a result, the computed monomial coefficients are generally highly inaccurate when the dimensionality of the Vandermonde matrix is not small. For this reason, other better-conditioned bases are often used for polynomial interpolation [32, 11]. On the other hand, it has long been observed that polynomial interpolation in the monomial basis produces highly accurate approximations for many smooth functions (see, for example, [17, 18]). This is because computing the monomial coefficients inaccurately does not imply that the resulting interpolating polynomial is necessarily a poor approximation, since it is the backward error ${\lVert V^{(N)}\widehat{a}^{(N)}-f^{(N)}\rVert}_{2}$ of the numerical solution $\widehat{a}^{(N)}$ to the Vandermonde system $V^{(N)}a^{(N)}=f^{(N)}$ that determines the approximation accuracy, and ${\lVert V^{(N)}\widehat{a}^{(N)}-f^{(N)}\rVert}_{2}$ can be small even when the condition number $\kappa(V^{(N)})$ is large. It is easy to show that ${\lVert V^{(N)}\widehat{a}^{(N)}-f^{(N)}\rVert}_{2}\lesssim u\cdot{\lVert a^{(N)}\rVert}_{2}$, where $u$ denotes machine epsilon, from which one can show that the monomial approximation error is bounded by the sum of the polynomial interpolation error and the extra error term $u\cdot{\lVert a^{(N)}\rVert}_{2}$. In this paper, we establish an upper bound for ${\lVert a^{(N)}\rVert}_{2}$ that depends on the polynomial interpolation errors $\{{\lVert F-P_{n}\rVert}_{L^{\infty}}\}_{n=0,1,\dots,N}$. Based on this bound, we find that this extra error term is generally smaller than the polynomial interpolation error, provided that the order of approximation is no larger than the maximum order allowed by the constraint $\kappa(V^{(N)})\lesssim\frac{1}{u}$ (see Section 2.1). This finding elucidates the unexpectedly good performance of polynomial interpolation using the monomial basis in practical applications. In the rare situations where the monomial basis results in a less accurate interpolant, we show that such instances can be easily identified and corrected a posteriori, leading to a robust algorithm for piecewise polynomial interpolation over simply connected compact regions in the complex plane using the monomial basis, with no extra error and with almost no extra cost (see Section 2.2). As a by-product of our analysis, we derive a tight upper bound for the condition number of any Vandermonde matrix (see Theorem 2.5), which includes several previously established results [14, 15] as special cases. Finally, we present applications where using the monomial basis for interpolation is particularly advantageous (see Section 4).
@@
While not discussed in this paper, our results can be easily generalized to higher dimensions and to polynomial approximants constructed using least-square fitting. In [30], we study bivariate polynomial interpolation in the monomial basis over a (possibly curved) triangle, and demonstrate that similarly accurate approximations can be constructed for orders up to 20, regardless of the triangle’s aspect ratio.